number of ways of selecting two numbers from 1 to 12 such that their sum is divisible by 3.

Grade:10

3 Answers

Pratham Ashish
14 years ago

Dear Naveen

for a no to be divisible by 3 , possibilities are 3,6,9,15,18,21

if 2 no. are x and y , den

eqns are

x + y = 3 ,x + y = 6, x + y = 9, x + y = 12 , x + y = 15 , x + y = 18 , x + y = 21

for 1st eqn x,y wud be 1 & 2

for 2nd set is (2,4),(1,5)

for 3rd (1,8),(2,7),(3,6),(4,5)

for 4th (1,11),(2,10),(3,9),(4,8),(5,7),(6,6)

for 5th (12,3),(11,4),(9,6),(10,5),(8,7)

for 6th (12,6),(11,7),(10,8)

for 7th (10,11)

just count the possibility and dat is ur ans

Harpreet Kaur Oberoi
14 years ago

12

saumendra aggarwal
14 years ago
48